OpenStudy (anonymous):

You start out with $20 and then spend money in a store where every item is $3. Y = -3x + 20. Using the graph you drew for the equation, predict the amount you will have left if you buy 4 items.

OpenStudy (p0sitr0n):

x = nb of items = 4 y = -3 *4 +20 y = -12 +20 y = 8$